# Privacy Guarantees in Posterior Sampling under Contamination

### v2

@@ C.3 Cauchy Regression Case
where $e_{i}\sim\text{Cauchy}(0,\gamma)$ is Cauchy random variable with unknown dispersion parameter $\gamma$. Assume that the covariates $\bm{W}_{i}$ satisfy $\|\bm{W}_{i}\|_{\infty}\leq 1$. Then the likelihood is given by
@@
$\left|\sqrt{f_{\theta,\gamma_{1}}(x|\bm{w})}-\sqrt{f_{\vartheta,\gamma_{2}}(x|\bm{w})}\right|\leq M^{(n)}_{\vartheta,\gamma_{2},r}(x,\bm{w})\|(\theta,\gamma_{1})-(\vartheta,\gamma_{2})\|_{\infty}$
@@
$M^{(n)}_{\vartheta,\gamma_{2},r}(x,\bm{w})=\frac{d}{\sqrt{\pi}}\left(1+\min_{\varphi\in B_{r}(\vartheta)}(x-\varphi^{\top}\bm{w})^{2}/R_{n}\right)^{-\frac{3}{2}}\\ \cdot\left[R_{n}^{\frac{3}{4}}(|x|+M_{n}\|\bm{w}\|_{2})\|\bm{w}\|_{2}+\frac{1}{4}R_{n}^{\frac{3}{4}}+R_{n}^{\frac{7}{4}}(|x|+M_{n}\|\bm{w}\|_{2})^{2}\right]$
@@
$\phi_{n}=n^{-1/q},\quad M_{n}=n^{k}$
for some $q>4$ and $k>0$ to be chosen for Assumption 3. It is not hard to check Lemma 7 holds for this likelihood and thus Assumption 2 is satisfied by applying Lemma 8 and Remark 14. Assumption 3 can be checked using the same procedure as in the Linear regression example which again gives $k>\tfrac{1}{4}$.
@@
$\epsilon_{n}\sim n^{-1/2q},\qquad\delta_{n}\sim\exp(-\tfrac{1}{2}c_{1}n^{1-2/q})$
@@
$\left|\sqrt{f_{\theta,\gamma_{1}}(x|\bm{w})}-\sqrt{f_{\vartheta,\gamma_{2}}(x|\bm{w})}\right|\leq M_{\vartheta,\gamma_{2},r}(x,\bm{w})\|(\theta,\gamma_{1})-(\vartheta,\gamma_{2})\|_{\infty}$
where function $M_{\vartheta,\gamma_{2},r}$ is given by
$M_{\vartheta,\gamma_{2},r}(x,\bm{w}):=\frac{d}{\sqrt{\pi}}\left(1+\min_{\varphi\in B_{r}(\vartheta)}(x-\varphi^{\top}\bm{w})^{2}/R_{n}\right)^{-\frac{3}{2}}\\ \cdot\left[R_{n}^{\frac{3}{4}}(|x|+M_{n}\|\bm{w}\|_{2})\|\bm{w}\|_{2}+\frac{1}{4}R_{n}^{\frac{3}{4}}+R_{n}^{\frac{7}{4}}(|x|+M_{n}\|\bm{w}\|_{2})^{2}\right]$
@@
$\nabla_{\theta}\sqrt{f_{\theta,\gamma}(x|\bm{w})}=\gamma^{\frac{1}{4}}\pi^{-\frac{1}{2}}\left(1+(x-\theta^{\top}\bm{w})^{2}/\gamma\right)^{-\frac{3}{2}}\frac{x-\theta^{\top}\bm{w}}{\gamma}\bm{w}$
$\nabla_{\gamma}\sqrt{f_{\theta,\gamma}(x|\bm{w})}=\frac{1}{4}\gamma^{-\frac{7}{4}}\pi^{-\frac{1}{2}}\left(1+(x-\theta^{\top}\bm{w})^{2}/\gamma\right)^{-\frac{3}{2}}\left[\gamma+3(x-\theta^{\top}\bm{w})^{2}\right]$
@@
$=$ $\,\xi^{\frac{1}{4}}\pi^{-\frac{1}{2}}\left(1+(x-\varphi^{\top}\bm{w})^{2}/\xi\right)^{-\frac{3}{2}}\frac{x-\varphi^{\top}\bm{w}}{\xi}(\theta-\vartheta)^{\top}\bm{x}$
@@
$\quad\cdot\left[R_{n}^{\frac{3}{4}}(|x|+M_{n}\|\bm{w}\|_{2})\|\bm{w}\|_{2}+\frac{1}{4}R_{n}^{\frac{3}{4}}+R_{n}^{\frac{7}{4}}(|x|+M_{n}\|\bm{w}\|_{2})^{2}\right]$
@@
$M_{\vartheta,\gamma_{2},r}(x,\bm{w}):=\frac{d}{\sqrt{\pi}}\left(1+\min_{\varphi\in B_{r}(\vartheta)}(x-\varphi^{\top}\bm{w})^{2}/R_{n}\right)^{-\frac{3}{2}}\\ \cdot\left[R_{n}^{\frac{3}{4}}(|x|+M_{n}\|\bm{w}\|_{2})\|\bm{w}\|_{2}+\frac{1}{4}R_{n}^{\frac{3}{4}}+R_{n}^{\frac{7}{4}}(|x|+M_{n}\|\bm{w}\|_{2})^{2}\right]$
@@
Again, the integral $\|M_{\vartheta,\gamma_{2},r}(x,\bm{w})\|_{2}^{2}$ can be evaluated separately in three intervals $\int_{-\infty}^{u_{-}}$,$\int_{u_{-}}^{u_{+}}$ and $\int_{u_{+}}^{\infty}$. By noting that $\max\{|u_{-}|,|u_{+}|\}\leq(\|\vartheta\|_{2}+r)\|\bm{w}\|_{2}\leq M_{n}\|\bm{w}\|_{2}$,
@@
Let $\|\bm{w}\|_{\infty}\leq 1$ with non-zero entries,
@@
Then we have for $(\theta,\gamma)$ close to $(\theta^{*},\gamma^{*})$, there exists positive constant $C$ independent of $(\theta,\gamma)$ such that
$h((\theta^{*},\gamma^{*}),(\theta,\gamma))\leq\phi_{n}\implies\|(\theta^{*},\gamma^{*})-(\theta,\gamma))\|_{2}\leq C\phi_{n}$
@@
$h^{2}((\theta^{*},\gamma^{*}),(\theta,\gamma))=\frac{1}{2}\int\left(\sqrt{\frac{\sqrt{\gamma^{*}}}{\pi(1+(x-\theta^{*\top}\bm{w})^{2}/\gamma^{*})}}-\sqrt{\frac{\sqrt{\gamma}}{\pi(1+(x-\theta^{\top}\bm{w})^{2}/\gamma)}}\right)^{2}\mathtt{d}y.$
@@
$\nabla_{\theta}\sqrt{f_{\theta,\gamma}(x|\bm{w})}=\gamma^{\frac{1}{4}}\pi^{-\frac{1}{2}}\left(1+(x-\theta^{\top}\bm{w})^{2}/\gamma\right)^{-\frac{3}{2}}\frac{x-\theta^{\top}\bm{w}}{\gamma}\bm{w}$
$\nabla_{\gamma}\sqrt{f_{\theta,\gamma}(x|\bm{w})}=\frac{1}{4}\gamma^{-\frac{7}{4}}\pi^{-\frac{1}{2}}\left(1+(x-\theta^{\top}\bm{w})^{2}/\gamma\right)^{-\frac{3}{2}}\left[\gamma+3(x-\theta^{\top}\bm{w})^{2}\right]$
@@
$=$ $\,\int\frac{\sqrt{\xi}}{\pi}\left(1+\frac{(x-\varphi^{\top}\bm{w})^{2}}{\xi}\right)^{-3}\left[\frac{x-\varphi^{\top}\bm{w}}{\xi}(\theta-\theta^{*})^{\top}\bm{x}+\frac{\gamma-\gamma^{*}}{4\xi}+\frac{3}{\xi^{2}}(x-\varphi^{\top}\bm{w})^{2}(\gamma-\gamma^{*})\right]^{2}\mathtt{d}x$
@@
The last step is due to the value of integral with respect to $\frac{x-\varphi^{\top}\bm{w}}{\xi}$ is only dependent on $\xi$ which is ‘close’ to $\gamma^{*}$ and thus has a finite lower and upper bound. Thus, since $\|\bm{w}\|_{\infty}\leq 1$ with non-zero entries,
$h^{2}((\theta^{*},\gamma^{*}),(\theta,\gamma))\succeq\|(\theta,\gamma)-(\theta^{*}-\gamma^{*})\|_{2}^{2}.$
@@ C.4 Quantile Regression Case
$Q_{\tau}(u):=\begin{dcases}(\tau-1)u,&u\leq 0\\ \tau u,&u>0.\end{dcases}$
If $\theta,\vartheta\in\Omega_{n}$ with $\|\theta-\vartheta\|_{\infty}\leq r$ then
$\left|\sqrt{f^{\tau}_{\theta}(x|\bm{w})}-\sqrt{f^{\tau}_{\vartheta}(x|\bm{w})}\right|\leq M_{\vartheta,r}(x,\bm{w})\|\theta-\vartheta\|_{\infty}$
where function $M_{\vartheta,r}$ is given by
@@
$\nabla_{\theta}\sqrt{f^{\tau}_{\theta}(x|\bm{w})}=\begin{dcases}-\sqrt{\tau(1-\tau)}(1-\tau)\exp\left(-\frac{\tau-1}{2}(x-\theta^{\top}\bm{w})\right)\bm{w},&y\leq\theta^{\top}\bm{w}\\ -\sqrt{\tau(1-\tau)}\tau\exp\left(-\frac{\tau}{2}(x-\theta^{\top}\bm{w})\right)\bm{w},&y>\theta^{\top}\bm{w}.\end{dcases}$
Using the same argument as in Lemma 20, recall the $\tilde{\nabla}$ operator from Definition 9 for the supremum of all direction derivatives. Let $\varphi=\theta+\alpha(\vartheta-\theta)$, then
@@
$\Omega_{n}:=\{\theta\in\Omega:\|\theta-\theta^{*}\|_{2}\leq R_{n}\}$
for some $R_{n}>0$. For $(\vartheta,\varsigma)\in\Omega_{n}$, let
@@
$\min_{\varphi\in B_{r}(\vartheta)}|x-\varphi^{\top}\bm{w}|=\begin{dcases}u_{-}-x,&x\leq u_{-}\\ 0,&u_{-}<x\leq u_{+}\\ x-u_{+},&x>u_{+}.\end{dcases}$
@@
Since $\|\bm{w}\|_{\infty}\leq 1$, the above expression does not depend on the radius of the space $\Omega_{n}$ and hence a constant order value. ∎
@@
where $\theta\in\Omega$. If $\|\bm{w}\|_{\infty}\leq 1$ with non-zero entries, then for $\theta$ close enough to $\theta^{*}$, there exists positive constant $C$ independent of $\theta$ such that
@@
$h^{2}(\theta,\theta^{*})=\frac{1}{2}\int\left(\sqrt{f_{\theta}^{\tau}(x|\bm{w})}-\sqrt{f_{\theta^{*}}^{\tau}(x|\bm{w})}\right)^{2}\mathtt{d}x$
@@
$\nabla_{\theta}\sqrt{f_{\theta}^{\tau}(x|\bm{w})}=\begin{dcases}-\sqrt{\tau(1-\tau)}\frac{1-\tau}{2}\exp\left(-\frac{\tau-1}{2}(x-\theta^{\top}\bm{w})\right)\bm{w},&x\leq\theta^{\top}\bm{w}\\ \sqrt{\tau(1-\tau)}\frac{\tau}{2}\exp\left(-\frac{\tau}{2}(x-\theta^{\top}\bm{w}\right)\bm{w},&x>\theta^{\top}\bm{w}.\end{dcases}$
@@
    $\sqrt{f_{\theta}^{\tau}(x|\bm{w})}-\sqrt{f_{\theta^{*}}^{\tau}(x|\bm{w})}=\sqrt{\tau(1-\tau)}\frac{\tau-1}{2}\exp\left(-\frac{\tau-1}{2}(x-\varphi_{1}^{\top}\bm{w})\right)\bm{x}^{\top}(\theta-\theta^{*}).$
@@
    $=$ $\,\frac{\tau(1-\tau)^{2}}{4}(\bm{w}^{\top}(\theta-\theta^{*}))^{2}\exp(-(\tau-1)(\theta^{*}-\varphi_{1})^{\top}\bm{x}).$
@@
Since the integrand in $h(\theta^{*},\theta)$ is non-negative, the integral is lower-bounded by the integral computed only on the regions described in the above two cases. When $\theta$ is close to $\theta^{*}$ and $\|\bm{w}\|_{\infty}\leq 1$ with non-zero entries,

### v3

@@ C.3 Cauchy Regression Case
where $e_{i}\sim\text{Cauchy}(0,\gamma)$ are Cauchy random variables with unknown dispersion parameter $\gamma$. Assume that the covariates $\bm{W}_{i}$ satisfy $\|\bm{W}_{i}\|_{\infty}\leq 1$. Then the likelihood is given by
@@
$\left|\sqrt{f_{\theta,\gamma_{1}}(x|\bm{w})}-\sqrt{f_{\vartheta,\gamma_{2}}(x|\bm{w})}\right|\leq M^{(n)}_{\vartheta,\gamma_{2},r}(x,\bm{w})\|(\theta,\gamma_{1})-(\vartheta,\gamma_{2})\|_{\infty},$
@@
$M^{(n)}_{\vartheta,\gamma_{2},r}(x,\bm{w})=\frac{d}{\sqrt{\pi}}\left(1+\min_{\varphi\in B_{r}(\vartheta)}(x-\varphi^{\top}\bm{w})^{2}/R_{n}\right)^{-\frac{3}{2}}\\ \cdot\left[R_{n}^{\frac{3}{4}}(|x|+M_{n}\|\bm{w}\|_{2})\|\bm{w}\|_{2}+\frac{1}{4}R_{n}^{\frac{3}{4}}+R_{n}^{\frac{7}{4}}(|x|+M_{n}\|\bm{w}\|_{2})^{2}\right],$
@@
$\phi_{n}=n^{-1/q},\quad M_{n}=n^{k},$
for some $q>4$ and $k>0$ to be chosen to satisfy Assumption 3. It is not hard to check that Lemma 6 holds for this likelihood and thus Assumption 2 is satisfied by applying Lemma 7 and Remark 12. Assumption 3 can be checked using the same procedure as in the linear regression example, which again gives $k>\tfrac{1}{4}$.
@@
$\epsilon_{n}\sim n^{-1/2q},\qquad\delta_{n}\sim\exp(-\tfrac{1}{2}c_{1}n^{1-2/q}),$
@@
$\left|\sqrt{f_{\theta,\gamma_{1}}(x|\bm{w})}-\sqrt{f_{\vartheta,\gamma_{2}}(x|\bm{w})}\right|\leq M_{\vartheta,\gamma_{2},r}(x,\bm{w})\|(\theta,\gamma_{1})-(\vartheta,\gamma_{2})\|_{\infty},$
where the function $M_{\vartheta,\gamma_{2},r}$ is given by
$M_{\vartheta,\gamma_{2},r}(x,\bm{w}):=\frac{d}{\sqrt{\pi}}\left(1+\min_{\varphi\in B_{r}(\vartheta)}(x-\varphi^{\top}\bm{w})^{2}/R_{n}\right)^{-\frac{3}{2}}\\ \cdot\left[R_{n}^{\frac{3}{4}}(|x|+M_{n}\|\bm{w}\|_{2})\|\bm{w}\|_{2}+\frac{1}{4}R_{n}^{\frac{3}{4}}+R_{n}^{\frac{7}{4}}(|x|+M_{n}\|\bm{w}\|_{2})^{2}\right],$
@@
$\begin{dcases}\nabla_{\theta}\sqrt{f_{\theta,\gamma}(x|\bm{w})}=\gamma^{\frac{1}{4}}\pi^{-\frac{1}{2}}\left(1+(x-\theta^{\top}\bm{w})^{2}/\gamma\right)^{-\frac{3}{2}}\frac{x-\theta^{\top}\bm{w}}{\gamma}\bm{w},\\ \nabla_{\gamma}\sqrt{f_{\theta,\gamma}(x|\bm{w})}=\frac{1}{4}\gamma^{-\frac{7}{4}}\pi^{-\frac{1}{2}}\left(1+(x-\theta^{\top}\bm{w})^{2}/\gamma\right)^{-\frac{3}{2}}\left[\gamma+3(x-\theta^{\top}\bm{w})^{2}\right].\end{dcases}$
@@
$=$ $\,\xi^{\frac{1}{4}}\pi^{-\frac{1}{2}}\left(1+(x-\varphi^{\top}\bm{w})^{2}/\xi\right)^{-\frac{3}{2}}\frac{x-\varphi^{\top}\bm{w}}{\xi}(\theta-\vartheta)^{\top}\bm{w}$
@@
$\quad\cdot\left[R_{n}^{\frac{3}{4}}(|x|+M_{n}\|\bm{w}\|_{2})\|\bm{w}\|_{2}+\frac{1}{4}R_{n}^{\frac{3}{4}}+R_{n}^{\frac{7}{4}}(|x|+M_{n}\|\bm{w}\|_{2})^{2}\right].$
@@
$M_{\vartheta,\gamma_{2},r}(x,\bm{w}):=\frac{d}{\sqrt{\pi}}\left(1+\min_{\varphi\in B_{r}(\vartheta)}(x-\varphi^{\top}\bm{w})^{2}/R_{n}\right)^{-\frac{3}{2}}\\ \cdot\left[R_{n}^{\frac{3}{4}}(|x|+M_{n}\|\bm{w}\|_{2})\|\bm{w}\|_{2}+\frac{1}{4}R_{n}^{\frac{3}{4}}+R_{n}^{\frac{7}{4}}(|x|+M_{n}\|\bm{w}\|_{2})^{2}\right],$
@@
Again, the integral $\|M_{\vartheta,\gamma_{2},r}(x,\bm{w})\|_{2}^{2}$ can be evaluated separately in three intervals $\int_{-\infty}^{u_{-}}$$\int_{u_{-}}^{u_{+}}$ and $\int_{u_{+}}^{\infty}$. By noting that $\max\{|u_{-}|,|u_{+}|\}\leq(\|\vartheta\|_{2}+r)\|\bm{w}\|_{2}\leq M_{n}\|\bm{w}\|_{2}$,
@@
Let $\|\bm{w}\|_{\infty}\leq 1$ with nonzero entries,
@@
Then we have for $(\theta,\gamma)$ close to $(\theta^{*},\gamma^{*})$, there exists a positive constant $C$ independent of $(\theta,\gamma)$ such that
$h((\theta^{*},\gamma^{*}),(\theta,\gamma))\leq\phi_{n}\implies\|(\theta^{*},\gamma^{*})-(\theta,\gamma))\|_{2}\leq C\phi_{n}.$
@@
$h^{2}((\theta^{*},\gamma^{*}),(\theta,\gamma))=\frac{1}{2}\int\left(\sqrt{\frac{\sqrt{\gamma^{*}}}{\pi(1+(x-\theta^{*\top}\bm{w})^{2}/\gamma^{*})}}-\sqrt{\frac{\sqrt{\gamma}}{\pi(1+(x-\theta^{\top}\bm{w})^{2}/\gamma)}}\right)^{2}\mathtt{d}x.$
@@
$\begin{dcases}\nabla_{\theta}\sqrt{f_{\theta,\gamma}(x|\bm{w})}=\gamma^{\frac{1}{4}}\pi^{-\frac{1}{2}}\left(1+(x-\theta^{\top}\bm{w})^{2}/\gamma\right)^{-\frac{3}{2}}\frac{x-\theta^{\top}\bm{w}}{\gamma}\bm{w},\\ \nabla_{\gamma}\sqrt{f_{\theta,\gamma}(x|\bm{w})}=\frac{1}{4}\gamma^{-\frac{7}{4}}\pi^{-\frac{1}{2}}\left(1+(x-\theta^{\top}\bm{w})^{2}/\gamma\right)^{-\frac{3}{2}}\left[\gamma+3(x-\theta^{\top}\bm{w})^{2}\right].\end{dcases}$
@@
$=$ $\,\int\frac{\sqrt{\xi}}{\pi}\left(1+\frac{(x-\varphi^{\top}\bm{w})^{2}}{\xi}\right)^{-3}\left[\frac{x-\varphi^{\top}\bm{w}}{\xi}(\theta-\theta^{*})^{\top}\bm{w}+\frac{\gamma-\gamma^{*}}{4\xi}+\frac{3}{\xi^{2}}(x-\varphi^{\top}\bm{w})^{2}(\gamma-\gamma^{*})\right]^{2}\mathtt{d}x$
@@
The last step is due to the value of integral with respect to $\frac{x-\varphi^{\top}\bm{w}}{\xi}$ is only dependent on $\xi$ which is ‘close’ to $\gamma^{*}$ and thus has a finite lower and upper bound. Thus, since $\|\bm{w}\|_{\infty}\leq 1$ with nonzero entries,
$h^{2}((\theta^{*},\gamma^{*}),(\theta,\gamma))\succeq\|(\theta,\gamma)-(\theta^{*},\gamma^{*})\|_{2}^{2}.$
@@ C.4 Quantile Regression Case
In quantile regression, one would like to estimate the $\tau$-th quantile, denoted $q_{\tau}(X_{i}|\bm{W}_{i})$ of the random variable $X_{i}$ conditioned on the covariates $\bm{W}_{i}$. As shown in Koenker and Bassett Jr (1978), solving the quantile regression problem is equivalent to solving the following optimisation problem:
$\min_{\theta}\sum_{i}Q_{\tau}(x_{i}-\theta^{\top}\bm{w}_{i}),\quad Q_{\tau}(u)=\frac{1}{2}(|u|+(2\tau-1)u).$
This is equivalent to maximising the likelihood
$\max_{\theta}\prod_{i}f_{\tau}(x_{i};\bm{w}_{i},\theta),$
where the probability density function $f_{\tau}(x;\bm{w}_{i},\theta)$ is given by
$f^{\tau}_{\theta}(x|\bm{w})=\tau(1-\tau)\exp\left(-Q_{\tau}(x-\theta^{\top}\bm{w})\right),$
which corresponds to the asymmetric Laplace distribution. Please refer to Yu and Moyeed (2001) for more details on Bayesian quantile regression. First, we check Assumption 1by using Lemma 2. We showed that, by Lemma 26, for any $\theta,\vartheta\in\Omega_{n}:=\{\theta\in\Omega:\|\theta-\theta^{*}\|_{2}\leq R_{n}\}$ such that $\|\theta-\vartheta\|_{\infty}\leq r$,
$\left|\sqrt{f^{\tau}_{\theta}(x|\bm{w})}-\sqrt{f^{\tau}_{\vartheta}(x|\bm{w})}\right|\leq M_{\vartheta,r}(x,\bm{w})\|\theta-\vartheta\|_{\infty},$
where $\tau^{*}=\max\{\tau,1-\tau\}$, $M_{n}=R_{n}+\|\theta^{*}\|_{2}$ and
Then by Lemma 27, assuming $\|\bm{w}\|_{\infty}\leq 1$,
Thus Assumption 1 can be satisfied with the choice $\phi_{n}=n^{-1/q},R_{n}=n^{k}$ for some $q>4$ and $k>0$ to be chosen for Assumption 3. It is not hard to check that the likelihood function satisfies the condition in Lemma 6 when $\theta$ is close enough to $\theta^{*}$, i.e., within $L_{2}$-distance of some constant $c_{r}$. Thus $S_{0,\alpha}(t_{n})$ contains an $L_{2}$-neighbourhood around $\theta^{*}$ with its radius proportional to $t_{n}$ up to a constant depending on $\theta^{*}$ and $c_{r}$. Assumption 2 is satisfied by applying Lemma 7 and Remark 12. For Assumption 3, we consider for large $R_{n}$ and apply Remark 13,
We need
$R_{n}^{2}=n^{2k}\geq 4c_{1}n\phi_{n}^{2}=4c_{1}n^{(q-2)/q},$
which can be satisfied by choosing k>(q2)/2qk>(q-2)/2q, for instance, $q=6$, $k>\tfrac{1}{6}$.
Finally, Lemma 28 also shows that $\phi_{n}$ and $\psi_{n}$ in Proposition 3 can be chosen to diminish at the same rate. Thus, we can choose $\psi_{n}\propto n^{-1/q}$ and $p_{n}=n^{-1/2q}$ which gives
$\epsilon_{n}\sim n^{-1/2q},\qquad\delta_{n}\sim\exp(-\tfrac{1}{2}c_{1}n^{1-2/q}),$
where $c_{1}$ is a fixed constant and $q>4$.
@@
$Q_{\tau}(u):=\begin{dcases}(\tau-1)u,&u\leq 0,\\ \tau u,&u>0.\end{dcases}$
If $\theta,\vartheta\in\Omega_{n}$ with $\|\theta-\vartheta\|_{\infty}\leq r$, then
$\left|\sqrt{f^{\tau}_{\theta}(x|\bm{w})}-\sqrt{f^{\tau}_{\vartheta}(x|\bm{w})}\right|\leq M_{\vartheta,r}(x,\bm{w})\|\theta-\vartheta\|_{\infty},$
where the function $M_{\vartheta,r}$ is given by
@@
$\nabla_{\theta}\sqrt{f^{\tau}_{\theta}(x|\bm{w})}=\begin{dcases}-\sqrt{\tau(1-\tau)}(1-\tau)\exp\left(-\frac{\tau-1}{2}(x-\theta^{\top}\bm{w})\right)\bm{w},&x\leq\theta^{\top}\bm{w},\\ -\sqrt{\tau(1-\tau)}\tau\exp\left(-\frac{\tau}{2}(x-\theta^{\top}\bm{w})\right)\bm{w},&x>\theta^{\top}\bm{w}.\end{dcases}$
Using the same argument as in Lemma 20, recall the $\tilde{\nabla}$ operator from Definition 9 for the supremum of all directional derivatives. Let $\varphi=\theta+\alpha(\vartheta-\theta)$, then
@@
$\Omega_{n}:=\{\theta\in\Omega:\|\theta-\theta^{*}\|_{2}\leq R_{n}\},$
for some $R_{n}>0$. For $\vartheta\in\Omega_{n}$, let
@@
$\min_{\varphi\in B_{r}(\vartheta)}|x-\varphi^{\top}\bm{w}|=\begin{dcases}u_{-}-x,&x\leq u_{-},\\ 0,&u_{-}<x\leq u_{+},\\ x-u_{+},&x>u_{+}.\end{dcases}$
@@
Since $\|\bm{w}\|_{\infty}\leq 1$, the above expression does not depend on the radius of the space $\Omega_{n}$ and is hence of constant order. ∎
@@
where $\theta\in\Omega$. If $\|\bm{w}\|_{\infty}\leq 1$ with nonzero entries, then for $\theta$ close enough to $\theta^{*}$, there exists a positive constant $C$ independent of $\theta$ such that
@@
$h^{2}(\theta,\theta^{*})=\frac{1}{2}\int\left(\sqrt{f_{\theta}^{\tau}(x|\bm{w})}-\sqrt{f_{\theta^{*}}^{\tau}(x|\bm{w})}\right)^{2}\mathtt{d}x,$
@@
$\nabla_{\theta}\sqrt{f_{\theta}^{\tau}(x|\bm{w})}=\begin{dcases}-\sqrt{\tau(1-\tau)}\frac{1-\tau}{2}\exp\left(-\frac{\tau-1}{2}(x-\theta^{\top}\bm{w})\right)\bm{w},&x\leq\theta^{\top}\bm{w},\\ \sqrt{\tau(1-\tau)}\frac{\tau}{2}\exp\left(-\frac{\tau}{2}(x-\theta^{\top}\bm{w})\right)\bm{w},&x>\theta^{\top}\bm{w}.\end{dcases}$
@@
    $\sqrt{f_{\theta}^{\tau}(x|\bm{w})}-\sqrt{f_{\theta^{*}}^{\tau}(x|\bm{w})}=\sqrt{\tau(1-\tau)}\frac{\tau-1}{2}\exp\left(-\frac{\tau-1}{2}(x-\varphi_{1}^{\top}\bm{w})\right)\bm{w}^{\top}(\theta-\theta^{*}).$
@@
    $=$ $\,\frac{\tau(1-\tau)^{2}}{4}(\bm{w}^{\top}(\theta-\theta^{*}))^{2}\exp(-(\tau-1)(\theta^{*}-\varphi_{1})^{\top}\bm{w}).$
@@
Since the integrand in $h(\theta^{*},\theta)$ is non-negative, the integral is lower-bounded by the integral computed only on the regions described in the above two cases. When $\theta$ is close to $\theta^{*}$ and $\|\bm{w}\|_{\infty}\leq 1$ with nonzero entries,